Heat is transferred to 6 kg of air which is initially at 100 kPa and B
0C until its temperature reaches 600
K and pressure reaches at 150 kPa. Determine the change in internal energy, the change in enthalpy,
the heat supplied and the work done in the following processes:
1. Constant Volume Process followed by Constant PressureProcess
2. Constant Pressure process followed by Constant VolumeProcess
Assume that air is an ideal gas. R=8.314 KJ/Kmol K, Cp=20.785 KJ/Kmol K and Cv =20.785
KJ/Kmol KMolecular weight of air =29. B = 42
1
Expert's answer
2021-09-30T10:20:35-0400



1. Constant Volume Process followed by Constant Pressure Process: At first, we have:

"n=\frac{6\,kg}{0.029\,kg/mol}=206.897\,mol\,of\,air
\\ \text{ }
\\ T_1=42\,°C=315\,K
\\ V_1=\frac{nRT_1}{P_1}
\\ V_1=\cfrac{(206.897\,mol)(8.314\frac{kPaL}{mol\,K})(315\,K)}{100\,kPa}
\\ V_1=5418.446\,L"


First, since V1=V2 (for the constant volume process):


"T_2=T_1\frac{P_2}{P_1}=(315\,K)(\frac{150\,kPa}{100\,kPa})=472.5\,K"


Then, we know that for an isochoric process "W=\intop pdV=0", this implies that the first law of thermodynamics is now "\Delta U = Q-W=Q=nC_v(T_2-T_1)" and for the enthalpy change "\Delta H = nC_p(T_2-T_1)".


"W=0\,J
\\ \Delta U =Q=(206.897\,mol)(20.785\frac{J}{mol\,K})(472.5-315)K
\\ \Delta U =Q= 677305.778\,J
\\ \Delta H=(206.897\,mol)(20.785\frac{J}{mol\,K})(472.5-315)K
\\ \Delta H= 677305.778\,J"


The second part is an isobaric process, where


"P_2=P_3=150\,kPa
\\ T_2=472.5\,K; T_3=600\,K
\\ V_2=5418.446\,L
\\ V_3=V_2\frac{T_3}{T_2}=(5418.446\,L)(\frac{600\,K}{472.5\,K})
\\ V_3=6880.566\,L"


Now we can find the rest of the terms knowing the same equations as before for "\Delta U" and "\Delta H":


"W=P_2(V_3-V_2)
\\ W=(150\,kPa)(6880.566-5418.446)L=219318\,J
\\ \Delta U =(206.897\,mol)(20.785\frac{J}{mol\,K})(600-472.5)K
\\ \Delta U =548295.153\,J
\\ Q=\Delta U + W = 767613.153\,J
\\ \Delta H=(206.897\,mol)(20.785\frac{J}{mol\,K})(600-472.5)K
\\ \Delta H= 548295.153\,J"


The conclusion for the first part is


"W=(0+219318)\,J=219318\,J=219.318\,kJ
\\ Q=(677305.778+767613.153)\,J=1444918.931\,J=1444.918\,kJ
\\ \Delta U=(677305.778+548295.153)\,J=1225600.931 \,J=1225.601\,kJ
\\ \Delta H=(677305.778+548295.153)\,J=1225600.931 \,J=1225.601\,kJ"


2. Constant Pressure process followed by Constant Volume Process


First, we know that


"n=\frac{6\,kg}{0.029\,kg/mol}=206.897\,mol\,of\,air
\\ \text{ }
\\ T_1=42\,°C=315\,K
\\ V_1=\frac{nRT_1}{P_1}
\\ V_1=\cfrac{(206.897\,mol)(8.314\frac{kPaL}{mol\,K})(315\,K)}{100\,kPa}
\\ V_1=5418.446\,L"


First, since P1=P2 (for the constant pressure process):


"V_2=V_1\frac{T_2}{T_1}=(5418.446\,L)(\frac{600\,K}{315\,K})=10320.849\,L"


Now, we know that for this process we can calculate the following:


"W=P_1(V_2-V_1)=(100\,kPa)(10320.849-5418.446)L=490240.3\,J
\\ \Delta U =(206.897\,mol)(20.785\frac{J}{mol\,K})(600-315)K
\\ \Delta U =1225600.931\,J
\\ Q=\Delta U + W = 1775841.231\,J
\\ \Delta H=(206.897\,mol)(20.785\frac{J}{mol\,K})(600-315)K
\\ \Delta H= 1225600.931\,J"

Now, the second part of this other process involves a process at constant volume, where


"V_2=V_3=10320.849\,L
\\ T_2=T_1\frac{P_2}{P_1}=(600\,K)(\frac{150\,kPa}{100\,kPa})=900\,K"


Then, since the volume is constant no work is done thus


"W=0\,J
\\ \Delta U =Q=(206.897\,mol)(20.785\frac{J}{mol\,K})(900-600)K
\\ \Delta U =Q= 1290106.244\,J
\\ \Delta H=(206.897\,mol)(20.785\frac{J}{mol\,K})(900-600)K
\\ \Delta H= 1290106.244\,J"


For the second part, we have the following


"W=(490240.3+0)\,J=490240.3\,J=490.240\,kJ
\\ Q=(1775841.231+1290106.244)\,J=3065947.475\,J=3065.947\,kJ
\\ \Delta U=(1225600.931+1290106.244)\,J=2515707.175\,J=2515.707\,kJ
\\ \Delta H=(1225600.931+1290106.244)\,J=2515707.175\,J=2515.707\,kJ"



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.